Skip to end of metadata
Go to start of metadata

Unscheduled:

  • October 24th: The red link to GSIX was currently down and also the link to gu-br2 went down because of a fiber problem which is fixed now.
  • October 17th: E-VPN service was unstable. Supplier upgraded the firmware on the switch and that has resolved the instability issue.
  • October 13th: A broken FPC was replaced in hb-r2
  • October 11th: BGP IPv4 sessions between NORDUnet and SUNET went down due to misconfiguration in NORDUnet NCS
  • October 7th: The green link to GSIX was down for few minutes. Customer did not have any maintenance. No RFO found.
  • October 4th: Problem in our optical network between hv-r2 and gbg7-r2.

Scheduled:

  • October 30th: Supplier performed replacement of fiber cable in Gävle on the red network.
  • October 26th: A service that is going between tug-r1 and mdh-r2 was recreated on our optical NMS.
  • October 16th: Fiber supplier performed repair/replacement of fiber cable, Morjärv-Kälvudden.
  • October 14th: A maintenance was carried out in one of the MDH routers
  • October 12th: We performed a planned storage maintenance on the Adobe Connect system for SUNET.
  • October 11th: A broken FPC card needs to be replaced in router in Göteborg.
  • October 7th: A new service was created between core routers in GBG7 and BOS
  • October 6th: SUNET relocated tne FPC card to new SUNET-C router.
  • October 3rd: Fiber optimization work was carried out in Fredhäll site on SUNET-C network.

Scope

Unscheduled

Scheduled

Total

Hardware

1 pcs 00:32

1 pcs 01:01

2 pcs 01:33

Link

1 pcs 00:04

3 pcs 06:06

4 pcs 06:10

None

3 pcs 15:22

5 pcs 1d 02:10

8 pcs 1d 17:32

Routing

0 pcs

0 pcs

0 pcs

Software

1 pcs 01:00

0 pcs

1 pcs 01:00

  • No labels